Solve the system by the substitution method. Determine whether the equations are independent, dependent, or inconsistent.
3y - x = 5
-x + 2y = -4

Answer by mananth(16946)

3y - x = 5
x= 3y-5
-x + 2y = -4
plug value of x
-(3y-5)+2y=-4
-3y+5+2y=-4
-y+5=-4
-5
-y=-4 -5
-y=-9
y=9
...
plug value of y in 1
3*(9)-x=5
27-x=5
x=22
...
m.ananth@hotmail.ca